The KL Partnership is owned equally by Kayla and Lisa. At the beginning of the year, Kayla’s basis is $20,000 and Lisa’s basis is $16,000. Partnership debt did not change from the beginning to the end of the tax year. KL reported the following income and expenses for the current tax year:
Sales revenue $150,000
Cost of sales 80,000
Distribution to Lisa 15,000
Depreciation expense 20,000
Utilities 14,000
Rent expense 18,000
Long-term capital gain 6,000
Payment to Mercy Hospital for Kayla’s medical expenses 12,000

Compute the answers to Chapter 21 problem 49.   When complete you will Prepare only the Form 1065 Schedule M-1 and M-2

Assume that the begining balance of the partner's capital account equals the beginning basis for the partner's.
Answer #1
A.The partnership's ordinary taxable income is determined as follows:
Item Value ($)
Sales revenue 1,50,000
Cost of sales -80,000
Depreciation expense -20,000
Utilities -14,000
Rent expense -18,000
(a) Total ordinary income 18,000
Separately stated items:
Long term capital gain $6,000
The distribution to Lisa is not deductible by the partnership. The payment to Mercy hospital for Kaylas expenses is treated as a distribution to Kayla in the amount of $ 12,000. Kayla may claim a deduction for medical expenses on her personal income tax return.
B. Kayla's basis in her partnership interest at the end of the tax year is:
Value ($)
Beginning basis 20,000
Share of partnership income =18,000/2 9,000
Share of separately stated income (Long-term capital gain) = 6,000/2 3,000
Drawings (payment for medical exp) -12,000
Ending basis in Interest=20+9+3-12 20,000
C. Lisa's basis in her partnership interest at the end of the tax year is:
Beginning basis 16,000
Share of partnership income =18,000/2 9,000
Share of separately stated items (Long-term capital gain) =6,000/2 3,000
Drawings -15,000
Ending basis in interest =16+9-3+15 13,000